OpenStudy (anonymous):

Thom worked 18 hours last week at the sporting goods store and earned $92.00. If he continues to earn the same hourly pay, how many additional hours must he work to earn another $138.00?

OpenStudy (anonymous):

We did this earlier this morning.

OpenStudy (anonymous):

i know but i cant remember ? :/

OpenStudy (anonymous):

money per week... 92 / 18 = $5.11 per week. 138.00/5.11 = 27 weeks.

OpenStudy (anonymous):

:/ i did that grr

OpenStudy (anonymous):

ok i did the next problem tell me if its right

OpenStudy (anonymous):

x/12.6 = 45/63 the answer i got was x=9 am i correct

OpenStudy (anonymous):

blexting am i correct

OpenStudy (anonymous):

Just a minute

OpenStudy (anonymous):

YEAH you are CORRECT.

OpenStudy (anonymous):

yes thank u finally im correct :)

OpenStudy (anonymous):

ok 5/x = 30/72 the answer i got is x=5 am i correct

OpenStudy (anonymous):

blexting

OpenStudy (anonymous):

Just a minute

OpenStudy (anonymous):

5 30 - = -- x 72 (5)(72) = 30x 360 = 30x divide by 30 12 = x

OpenStudy (anonymous):

ok i got two different answers for this one 54/x = 9/7 i got x= 42 and x= 7

OpenStudy (anonymous):

54 9 -- = -- x 7 54(7) = 9x 378 = 9x divide by 9 42 = x

OpenStudy (anonymous):

ok so was i right

OpenStudy (anonymous):

x = 42 is the only answer.

OpenStudy (anonymous):

ok sorry i just suck at math it to hard